# Enabling decision support over confidential data

### v1

@@ V Formal Analysis
In this section, we formalize SPARTA as a protocol to discuss its security guarantees. In Sect. V-A we present the necessary tools for the discussion, namely the notation, the Random Oracle and finally the Authenticated Encryption. Then, in Sect. V-B we introduce the properties that we require for our protocol (see Sect. V-B1). We then present a thread model of our solution (see Sect. V-B2), and we finally conclude with the Theorem and its proof in Sect. V-B3.
V-A Useful Tools
@@
(1) Notation.Let $\mathbb{N}$ be the set of all natural numbers. We denote the security parameter by $\lambda\in\mathbb{N}$. Every algorithm takes the security parameter $\lambda$ as input (in unary, i.e., $1^{\lambda}$). When an algorithm has multiple inputs, $1^{\lambda}$ is typically omitted. A function is called negligible if, for every positive polynomial and all sufficiently large security parameter $\lambda$, the function grows at most as the inverse of the polynomial in $\lambda$. (2) Random Oracle.In the Random Oracle Model (ROM), all the parties have access to a truly random hash function. In particular, when a value $v$ is given as an input from a party to the Random Oracle (RO), the latter samples a random answer $r$, stores the pair $(v,r)$, and outputs $r$ to the party. If the RO is queried on the same value $v$ multiple times, the same answer $r$ is output. It is heuristically shown that cryptographic hash functions such as $\mathsf{SHA\text{-}256}$ behave as a random oracle. (3) Authenticated Encryption.A Symmetric Key Encryption scheme with key space $\mathcal{K}$ is composed of the following algorithms: (i) $\mathsf{Gen}(1^{\lambda})$: The randomized key generation algorithm outputs a key $k$ in the key space $\mathcal{K}$. (ii) $\mathsf{Enc}(k,m)$: The randomized encryption algorithm, on input the key $k$ and a message $m$, outputs a ciphertext $c$. (iii) $\mathsf{Dec}(k,c)$: The decryption algorithm, on input the key $k$ and a ciphertext $c$, outputs a message $m$, or $\bot$ if decryption fails. An authenticated encryption scheme should satisfy the following properties: (i) Correctness:For each message $m$ and each $k\in\mathcal{K}$, $\Pr[\mathsf{Dec}(k,\mathsf{Enc}(k,m))=m]=1$. (ii) Unforgeability:The probability that a PPT adversary $\mathcal{A}$ wins the following experiment (i.e., outputs 1) is negligible: (a) Obtain $k$ from $\mathsf{Gen}(1^{\lambda})$. (b) The adversary, after making at most polynomially many queries to the encryption oracle $\mathsf{Enc}(k,\cdot)$, outputs a ciphertext $c$. (c) Let $m=\mathsf{Dec}(k,c)$. Output 1 if $m\neq\bot$ and $c$ was not queried to the encryption oracle by $\mathcal{A}$. Output 0 otherwise. (iii) CCA-Security:The probability that a PPT adversary $\mathcal{A}$ wins the following experiment (i.e., outputs 1) is negligibly close to $1/2$: (a) Obtain $k$ from $\mathsf{Gen}(1^{\lambda})$. (b) The adversary, with oracle access to $\mathsf{Enc}(k,\cdot)$ and $\mathsf{Dec}(k,\cdot)$, outputs messages $m_{0},m_{1}$. (c) Choose a uniform bit $b\in\{0,1\}$, compute ciphertext $c=\mathsf{Enc}(k,m_{b})$, and send $c$ to $\mathcal{A}$. (d) The adversary, with oracle access to $\mathsf{Enc}(k,\cdot)$ and restricted oracle access to $\mathsf{Dec}(k,\cdot)$ (excluding $c$), outputs a bit $b^{\prime}$. (e) The experiment outputs 11 if $b^{\prime}=b$, else 0.
V-B Formal Analysis
V-B1 Desired Properties
@@
(1) Correctness.For any function $f$ and any set of messages $M$ such that $f(M)=y$, if $M$ is provided to the CCU for function $f$ by the Data Provider obtaining ciphertexts $C$, and subsequently, ciphertexts $C$ are provided by the Decider back to the CCU obtaining output $y^{\prime}$, then $y^{\prime}=f(M)=y$. (2) Data Privacy.A PPT adversary $\mathcal{A}$ wins the following experiment with probability negligibly close to $1/2$: (i) $\mathcal{A}$chooses a function $f$. $\mathcal{A}$ has oracle access to the CCU acting as Data Provider (sending arbitrary sets of messages $M$ and receiving corresponding ciphertexts $C$) and oracle access to the CCU acting as Decider (sending arbitrary sets of ciphertexts $C$ and receiving back $f(M)$ if $M$ correctly decrypts from $C$). Then, $\mathcal{A}$ chooses two sets of messages $M_{0},M_{1}$ such that $f(M^{\prime}_{0})=f(M^{\prime}_{1})$ where $M^{\prime}_{b}$ can be any combination of messages in $M_{b}$. (ii) Choose a random bit $b$, produce ciphertexts $C_{b}$ of messages $M_{b}$, and send $C_{b}$ to $\mathcal{A}$. (iii) $\mathcal{A}$has oracle access to the CCU as a Data Provider to send arbitrary sets of messages $M$, and obtain the corresponding ciphertexts $C$ back. It has further oracle access to the CCU as a Decider by sending arbitrary ciphertexts $C$ and obtaining back $f(M)$ if $M$ correctly decrypts $C$. Finally, outputs a bit $b^{\prime}$. (iv) If $b^{\prime}=b$, output 1; otherwise, output 0. (3) Data Integrity.A PPT adversary $\mathcal{A}$ wins the following experiment with negligible probability: (i) $\mathcal{A}$, acting as Data Provider, has oracle access to the CCU submitting arbitrary message sets $M$ and receiving ciphertexts $C$. $\mathcal{A}$ outputs ciphertexts $C^{*}$. (ii) Decrypting ciphertexts $C^{*}$ yields messages $M^{*}$. Output 1 if no message in $M^{*}$ equals $\bot$ and no subset of $M^{*}$ was previously queried; else, output 0. (4) Forward/Backward Secrecy/Integrity.If all ciphertext keys at time $t$ are leaked, ciphertexts encrypted at any different time $t^{\prime}$ remain secure, and their underlying messages cannot be decrypted or forged.
V-B2 Threat Model
We assume the Data Provider and Certification Authority are honest, the Decider is semi-honest, [^8] We could assume a dishonest Decider by adding signatures on the decision and data produced by the CCU, but this is unnecessary in our setting. and secure authenticated channels can be established between authorized parties (e.g., CA, Decider) and the CCU. We model the CCU as a fully trusted third party computing decisions through functions $\{f_{j}\}$ using a hardwired secret seed $s$.
V-B3 The Protocol
Let $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$ be an authenticated encryption scheme with key space $\mathcal{K}$ and $H:\{0,1\}^{*}\rightarrow\mathcal{K}$ be a cryptographic hash function modeled as a random oracle. The protocol is as follows:
@@ V-B3 The Protocol
We are now ready to prove the following informal theorem.
@@
Recall that the CCU is modeled as a fully trusted third party. Hence, the fixed functions $\{f_{j}\}$ cannot be modified, internal states remain secret, and outputs are always correctly computed with respect to these functions.
Correctness is ensured by the correctness property of the authenticated encryption scheme $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$ and by the fact that the CCU’s functions are fixed and unmodifiable.
Data Privacy is guaranteed since the internal CCU state (i.e., the seed $s$) remains secret and the derived encryption keys $k_{i}=H(s,T_{i})$ appear indistinguishable from random due to the Random Oracle assumption. Thus, Data Privacy reduces directly to the CCA-security of $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$.
Data Integrity follows from the Unforgeability property of the authenticated encryption scheme $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$, combined with the trusted execution guarantees provided by the CCU.
Forward/Backward Secrecy and Integrity result directly from the random-oracle-based key derivation: given a leaked key $k=H(s,T)$, the one-way property of the Random Oracle prevents recovery of the seed $s$. Therefore, keys corresponding to ciphertexts encrypted at other timestamps $T^{\prime}\neq T$ remain computationally infeasible to derive. ∎
@@
VI Implementation and evaluation
We implemented SPARTA in Go, using the EGo framework to deploy trusted applications within an Intel SGX TEE. [^9] EGo: edgeless.systems/products/ego; Intel SGX: sgx101.gitbook.io/sgx101. Accessed 2025-06-17 We realized the Request Manager as a Redis queue. To evaluate FEEL filtering statements and access policies we resort to govaluate. For the evaluation of aggregation functions we rely on Expr. [^10] Redis queue: github.com/rq/rq; govaluate: github.com/Knetic/govaluate; Expr: github.com/expr-lang/expr. Accessed 2025-06-17 Our tests were conducted on a TEE hosted by an Intel Xeon Gold 5415+ CPU.
The objective of our experimental evaluation is to demonstrate the feasibility of our solution and assess its performance in multiple application scenarios. We gauge not only runtime but also computation memory in light of the limits imposed by TEEs (see Sect. II). In particular, we aim to answer the following core questions: 1) Does SPARTA scale efficiently with increasing input data size and decision table complexity? 2) Can SPARTA operate end-to-end on realistic input data and policy configurations across multiple domains? 3) What is the computational overhead introduced by encryption and the use of a TEE? To answer these questions and test distinct features of our system, we used the following three datasets.
@@ VI Implementation and evaluation
We generated this dataset based on our vaccine campaign use case (see Sect. III), incorporating realistic access control policies and data aggregation requirements relevant to public health scenarios. This synthetic dataset is essential to assess the practical feasibility of SPARTA and, due to the numerous data fields, to measure the memory savings achieved through our aggregation mechanism described in the WS step.
@@
This dataset originates from the February 2021 Decision Management Community challenge Determine Medical Service Coverage and contains over 16 00016\,000 healthcare-related records. Its large size makes it suitable for testing the scalability of SPARTA under realistic conditions, including measuring execution time with large input and evaluating performance when varying the complexity and size of the decision tables. The nature of the data (medical services and patient records) aligns well with scenarios requiring strong data confidentiality, allowing us to test encryption strategies and directly compare SPARTA’s performance inside a TEE versus a traditional server execution.
@@
VI-1 Running example tests
We assess the practical feasibility of SPARTA and quantify the memory savings achieved by the data aggregation mechanism (see the WS phase) using the synthetic VAX dataset, which simulates our motivating scenario of a nationwide vaccine campaign. Following the setup in Sect. III, we generated realistic input data for patients, vaccination centers, and carriers (Tab. I), aligning each record with the decision logic defined in the corresponding DMN tables (see Tab. IV for our running example, and the appendix for the other two decision functions). Each rule in the decision logic is matched by at least one input, and we manually verify the correctness of the computed decisions. To ensure reproducibility, all data generators, inputs, and outputs are publicly available in our repository.111111anonymous.4open.science/r/SPARTA
@@ VI-1 Running example tests
Figure 2: Execution time with increasing records and columns
VI-2 Determine Medical Service Coverage tests
@@
(a) Execution time with increasing number of records and columns
Figure 3 reports the execution time of the decision function (List. 2) under increasing workload. In Fig. 3(a), we fix the number of columns in the DMN to 771414, and $28$, and progressively increase the number of input records from $1000$ to 16 00016\,000. In Fig. 3(b), we fix the number of records to 1000100080008000, and 16 00016\,000, and vary the number of columns from 1 to 28. In both configurations, we observe a linear increase in execution time as the number of records and columns grows. The linear scalability shows that our solution scaled predictably with increasing data size and complexity. Such a characteristic is essential for real-world deployments where decision tables and input data may grow significantly.
@@
(b) Execution time with increasing number of rules
To assess the impact of the complexity of the decision on SPARTA’s runtime, we perform a test by incrementally increasing the number of rules involved. Figure 5 shows the result with a fixed number of records (16 00016\,000) and columns ($28$) while varying the number of rules from $100$ up to $1500$. The observed trend is linear, thus confirming a good scalability of our system also with regard to the number of decision rules.
(c) Execution time with function aggregation results as additional inputs
As described in the SPEC step, the Policymaker may define aggregation functions to be computed prior to the decision process, with the resulting values used as additional inputs. To test the overhead introduced by this feature, we extend the original $7$ columns by adding up to $21$ aggregation functions for a total of $28$ columns as input. From the eighth column onward, each column corresponds to one the aggregation output. Figure 5 shows the result with fixed 16 00016\,000 records. The figure shows that the decryption and aggregation time introduce a small overhead to the computation. The total time is comparable to the decision time. This result entails that the decryption and aggregation time are features that can be included in the system adding a small time overhead.
@@
Figure 4: Execution time with increasing number of rules
@@
Figure 5: Execution time with prior aggregation
(d) Memory usage
We measure the memory usage within the TEE (the enclave memory) during full execution runs. Figure 6 shows the memory Random Access Memory (RAM) usage for two different scenario: Figure 6(a) presents the results with prior aggregation, while Fig. 6(b) without. Each run begins with the initialization of the CCU and ends when the decision process terminates. The two images demonstrate a reasonable memory footprint throughout the computation with an expected spike in correspondence with data decryption.
@@
(e) Encryption overhead
To assess the overhead, we conducted experiments that account not only for the decision-making process itself but also for the data decryption phase. This step is essential in our system, as data must be decrypted after retrieval from IPFS and before decision execution (see the DECM step). We design two test configurations to evaluate the impact of different encryption strategies: light encryption, which serves as a baseline for comparison, and heavy encryption. Recalling the WS-ii step, in the heavy configuration, the CCU generates a different random number for each record, so it encrypts each such record with a different symmetric key (as per Tabs. II and II). For the light encryption variant, we modified the process described in WS-ii to use a single symmetric key (namely, we used the same random number for generating the symmetric key) for all entries. Indeed, this setup is strictly for experimental purposes. In practice, such an approach is discouraged, as compromising the single key would allow an attacker to decrypt the entire dataset rather than a single record. Figure 7 shows the execution time of the two encryption solutions. Figure 7(a) displays the execution time of light and heavy encryption strategies by varying the number of records with a fixed amount of columns ($28$), while in Fig. 7(b) we increase the number of columns with a fixed amount of users (16 00016\,000). The results highlight that the heavy encryption approach incurs a modest but reasonable overhead ($\approx\qty$80ms at most) compared to the light variant. However, this performance trade-off is justified because per-record encryption ensures significantly higher data confidentiality and resilience against key compromise.
@@
(f) TEE vs Server
To quantify the overhead introduced by using both TEE and strong cryptographic guarantees, we compare our decision support system running inside the TEE with a standard server-based execution, inspired by [47, 48]. Specifically, we modify our system to run outside the TEE but on the same machine. Figure 8 shows the comparison result: Figure 8(a) presents the execution time with increasing number of records and fixed columns, while in Fig. 8(b) we vary the number of columns with a fixed number of records (16 00016\,000). The results shows that the additional overhead is limited and primarily due to the decryption phase required by the TEE. In contrast, the server execution assume direct access to plaintext data within a trusted environment. Overall, our system provides strong security guarantees, including confidentiality for both data and decision logic, at a minimal performance cost and without compromising efficiency.
@@
VI-3 Card Approval tests
@@
Figure 9: Execution time with VAX, MED, and CARD datasets
@@
Figure 9 summarizes execution times across all evaluated datasets. Notably, the MED dataset incurs higher decision evaluation time due to its larger rule set (see Sect. VI-2). However, decryption time remains consistent, underscoring SPARTA’s ability to maintain stable performance across domains with varying logic complexity and input scale.
VII Limitations and Future Work
Our solution presents some limitations. Firstly, our current design involves a single TEE. Even though SPARTA allows for the replication of CCUs, a load balancing strategy is not yet implemented. Such an extension would require non-trivial communication and session management adjustments, but would allow for a large-scale adoption.
The memory limits of a TEE (see Sect. II) presents a scalability challenge when it comes to make decisions over a large volume of records. Solutions to this problem can involve integrating Oblivious RAM (ORAM) techniques [49, 50, 51, 52] or the execution of decisions in batched segments that can circumvent those memory constraints.
We developed our solution for Intel SGX TEEs. Other solutions exist (e.g., AMD SEV, Arm CCA) and studying their employment is an interesting endeavor. Also, we will consider the integration of Intel Trust Domain Extensions (TDX). Our implementation does not support the entire FEEL language to write the filtering expressions and the aggregation functions. In particular, we currently back only one operation over the filtered data, and support only the standard aggregation functions (e.g., meansummaxmin). To further increase the expressive power of the decision making processes, we envision the adoption of TEE-based smart contracts for automated decision-making [53, 54, 55, 56, 57].
Finally, we conducted experiments with public benchmarks showing good performance and scalability of our approach with low overhead. Nevertheless, an on-field use of our system in real-world settings is paramount for validation purposes.
VIII Related Work

### v2

@@ 5 Formal Analysis
In this section, we formalize SPARTA as a protocol to discuss its security guarantees. In Sect. 5.1 we present the necessary tools for the discussion, namely the notation, the Random Oracle and the Authenticated Encryption. Then, in Sect. 5.2 we introduce the properties that we require for our protocol (see Sect. 5.2.1). We then present a thread model of our solution (see Sect. 5.2.2), and we finally conclude with the Theorem and its proof in Sect. 5.2.3.
5.1 Useful Tools
Notation.
Let $\mathbb{N}$ be the set of all natural numbers. We denote the security parameter by $\lambda\in\mathbb{N}$. Every algorithm takes the security parameter $\lambda$ as input (in unary, i.e., $1^{\lambda}$). When an algorithm has multiple inputs, $1^{\lambda}$ is typically omitted. A function is called negligible if, for every positive polynomial and all sufficiently large security parameter $\lambda$, the function grows at most as the inverse of the polynomial in $\lambda$.
Random Oracle.
In the Random Oracle Model (ROM), all the parties have access to a truly random hash function. In particular, when a value $v$ is given as an input from a party to the Random Oracle (RO), the latter samples a random answer $r$, stores the pair $(v,r)$, and outputs $r$ to the party. If the RO is queried on the same value $v$ multiple times, the same answer $r$ is output. It is heuristically shown that cryptographic hash functions such as SHA-256 behave as a RO.
Authenticated Encryption.
@@ 5.1 Useful Tools
A Symmetric Key Encryption scheme with key space $\mathcal{K}$ is composed of the following algorithms: (i) $\mathsf{Gen}(1^{\lambda})$: The randomized key generation algorithm outputs a key $k$ in the key space $\mathcal{K}$. (ii) $\mathsf{Enc}(k,m)$: The randomized encryption algorithm, on input the key $k$ and a message $m$, outputs a ciphertext $c$. (iii) $\mathsf{Dec}(k,c)$: The decryption algorithm, on input the key $k$ and a ciphertext $c$, outputs a message $m$, or $\bot$ if decryption fails. An authenticated encryption scheme should satisfy the following properties: (i) Correctness:For each message $m$ and each $k\in\mathcal{K}$, $\Pr[\mathsf{Dec}(k,\mathsf{Enc}(k,m))=m]=1$. (ii) Unforgeability:The probability that a PPT adversary $\mathcal{A}$ wins the following experiment (i.e., outputs 1) is negligible: (a) Obtain $k$ from $\mathsf{Gen}(1^{\lambda})$. (b) The adversary, after making at most polynomially many queries to the encryption oracle $\mathsf{Enc}(k,\cdot)$, outputs a ciphertext $c$. (c) Let $m=\mathsf{Dec}(k,c)$. Output 1 if $m\neq\bot$ and $c$ was not queried to the encryption oracle by $\mathcal{A}$. Output 0 otherwise. (iii) CCA-Security:The probability that a PPT adversary $\mathcal{A}$ wins the following experiment (i.e., outputs 1) is negligibly close to $1/2$: (a) Obtain $k$ from $\mathsf{Gen}(1^{\lambda})$. (b) The adversary, with oracle access to $\mathsf{Enc}(k,\cdot)$ and $\mathsf{Dec}(k,\cdot)$, outputs messages $m_{0},m_{1}$. (c) Choose a uniform bit $b\in\{0,1\}$, compute ciphertext $c=\mathsf{Enc}(k,m_{b})$, and send $c$ to $\mathcal{A}$. (d) The adversary, with oracle access to $\mathsf{Enc}(k,\cdot)$ and restricted oracle access to $\mathsf{Dec}(k,\cdot)$ (excluding $c$), outputs a bit $b^{\prime}$. (e) The experiment outputs $1$ if $b^{\prime}=b$, else 0.
5.2 Formal Analysis
5.2.1 Desired Properties
@@
Correctness.
For any function $f$ and any set of messages $M$ such that $f(M)=y$, if $M$ is provided to the CCU for function $f$ by the Data Provider obtaining ciphertexts $C$, and subsequently, ciphertexts $C$ are provided by the Decider back to the CCU obtaining output $y^{\prime}$, then $y^{\prime}=f(M)=y$.
Data Privacy.
A PPT adversary $\mathcal{A}$ wins the following experiment with probability negligibly close to $1/2$: (i) $\mathcal{A}$chooses a function $f$. $\mathcal{A}$ has oracle access to the CCU acting as Data Provider (sending arbitrary sets of messages $M$ and receiving corresponding ciphertexts $C$) and oracle access to the CCU acting as Decider (sending arbitrary sets of ciphertexts $C$ and receiving back $f(M)$ if $M$ correctly decrypts from $C$). Then, $\mathcal{A}$ chooses two sets of messages $M_{0},M_{1}$ such that $f(M^{\prime}_{0})=f(M^{\prime}_{1})$ where $M^{\prime}_{b}$ can be any combination of messages in $M_{b}$. (ii) Choose a random bit $b$, produce ciphertexts $C_{b}$ of messages $M_{b}$, and send $C_{b}$ to $\mathcal{A}$. (iii) $\mathcal{A}$has oracle access to the CCU as a Data Provider to send arbitrary sets of messages $M$, and obtain the corresponding ciphertexts $C$ back. It has further oracle access to the CCU as a Decider by sending arbitrary ciphertexts $C$ and obtaining back $f(M)$ if $M$ correctly decrypts $C$. Finally, outputs a bit $b^{\prime}$. (iv) If $b^{\prime}=b$, output 1; otherwise, output 0.
Data Integrity.
A PPT adversary $\mathcal{A}$ wins the following experiment with negligible probability: (i) $\mathcal{A}$, acting as Data Provider, has oracle access to the CCU submitting arbitrary message sets $M$ and receiving ciphertexts $C$. $\mathcal{A}$ outputs ciphertexts $C^{*}$. (ii) Decrypting ciphertexts $C^{*}$ yields messages $M^{*}$. Output 1 if no message in $M^{*}$ equals $\bot$ and no subset of $M^{*}$ was previously queried; else, output 0.
Forward/Backward Secrecy/Integrity.
If all ciphertext keys at time $t$ are leaked, ciphertexts encrypted at any different time $t^{\prime}$ remain secure, and their underlying messages cannot be decrypted or forged.
5.2.2 Threat Model
We assume that Data Provider and CA are honest, the Decider is dishonest and secure authenticated channels can be established between authorized parties (e.g., CA, Decider) and the CCU. We model the CCU as a fully trusted third party computing decisions through functions $\{f_{j}\}$ using a hardwired secret seed $s$.
5.2.3 The Protocol
Let $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$ be an authenticated encryption scheme with key space $\mathcal{K}$ and $H:\{0,1\}^{*}\rightarrow\mathcal{K}$ be a cryptographic hash function modeled as a RO. The protocol is as follows:
@@ 5.2.3 The Protocol
No we can prove the following informal theorem.
@@
Recall that the CCU is modeled as a fully trusted third party. Hence, the fixed functions $\{f_{j}\}$ cannot be modified, internal states remain secret, and outputs are always correctly computed with respect to these functions. Correctness is ensured by the correctness property of the authenticated encryption scheme $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$ and by the fact that the CCU’s functions are fixed and unmodifiable. Data Privacy is guaranteed since the internal CCU state (i.e., the seed $s$) remains secret and the derived encryption keys $k_{i}=H(s,T_{i})$ appear indistinguishable from random due to the RO assumption. Thus, Data Privacy reduces directly to the CCA-security of $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$. Data Integrity follows from the Unforgeability property of the authenticated encryption scheme $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$, combined with the trusted execution guarantees provided by the CCU. Forward/Backward Secrecy and Integrity results directly from the RO-based key derivation. Given a leaked key $k=H(s,T)$, the one-way property of the RO prevents recovery of the seed $s$. Therefore, keys corresponding to ciphertexts encrypted at other timestamps $T^{\prime}\neq T$ remain computationally infeasible to derive. ∎
@@
6 Implementation and evaluation
We implemented SPARTA in Go, using the EGo framework to deploy trusted applications within an Intel SGX TEE. [^8] EGo: edgeless.systems/products/ego; Intel SGX: sgx101.gitbook.io/sgx101. Accessed 2025-06-17 We realized the Request Manager as a Redis queue. To evaluate FEEL filtering statements and access policies we resort to govaluate. We evaluated the aggregation functions with Expr. [^9] Redis queue: github.com/rq/rq; govaluate: github.com/Knetic/govaluate; Expr: github.com/expr-lang/expr. Accessed 2025-06-17 We conducted the tests on a TEE hosted by an Intel Xeon Gold 5415+ CPU.
The objective of our experimental evaluation is to demonstrate the feasibility of our solution and assess its performance in multiple application scenarios. We gauge not only runtime but also computation memory in light of the limits imposed by TEEs (see Sect. 2). In particular, we aim to answer the following core questions: 1) Does SPARTA scale efficiently with increasing input data size and decision table complexity? 2) Can SPARTA operate end-to-end on realistic input data and policy configurations across multiple domains? 3) What is the computational overhead introduced by encryption and the use of a TEE? To answer these questions and test distinct features of our system, we used the following three datasets.
@@ 6 Implementation and evaluation
We generated this dataset based on our vaccine campaign use case (see Sect. 3), incorporating realistic access control policies and data aggregation requirements relevant to public health scenarios. With this synthetic dataset we assess the practical feasibility of SPARTA and, due to the numerous data fields, to measure the memory savings achieved through our aggregation mechanism described in the WS step.
@@
This dataset originates from the February 2021 Decision Management Community challenge Determine Medical Service Coverage and contains over 16 00016\,000 healthcare-related records. Its large size allows for testing the scalability of SPARTA under realistic conditions, including measuring execution time with large input and evaluating performance when varying the complexity and size of the decision tables. The nature of the data (medical services and patient records) aligns well with scenarios requiring strong data confidentiality, allowing us to test encryption strategies and directly compare SPARTA’s performance inside a TEE versus a traditional server execution.
@@
Running example tests
@@
Figure 2: Memory savings for each decision
Determine Medical Service Coverage tests
@@
(a) Execution time with increasing number of records and columns Figure 3reports the execution time of the decision function (List. 2) under increasing workload. In Fig. 3(a), we fix the number of columns in the DMN to 77, $14$, and $28$, and progressively increase the number of input records from $1000$ to 16 00016\,000. In Fig. 3(b), we fix the number of records to $1000$, $8000$, and 16 00016\,000, and vary the number of columns from 1 to 28. In both configurations, we observe a linear increase in execution time as the number of records and columns grows. The linear scalability shows that our solution scaled predictably with increasing data size and complexity. Such a characteristic is essential for real-world deployments where decision tables and input data may grow significantly.
Refer to caption (a)
Refer to caption (b)
@@
(b) Execution time with increasing number of rules To assess the impact of the complexity of the decision on SPARTA’s runtime, we perform a test by incrementally increasing the number of rules involved. Figure 7 shows the result with a fixed number of records (16 00016\,000) and columns ($28$) while varying the number of rules from $100$ up to $1500$. The observed trend is linear, thus confirming a good scalability of our system also with regard to the number of decision rules. (c) Execution time with function aggregation results as additional inputs As described in the SPEC step, the Policymaker may define aggregation functions to be computed prior to the decision process, with the resulting values used as additional inputs. To test the overhead introduced by this feature, we extend the original 77 columns by adding up to 2121 aggregation functions for a total of $28$ columns as input. From the eighth column onward, each column corresponds to one the aggregation output. Figure 7 shows the result with fixed 16 00016\,000 records. The figure shows that the decryption and aggregation time introduce a small overhead to the computation. The total time is comparable to the decision time. This result entails that the decryption and aggregation are features that can be included in the system adding a small time overhead.
@@
[…]
(d) Memory usage We measure the memory usage within the TEE (the enclave memory) during full execution runs. Figure 8 shows the memory Random Access Memory (RAM) usage for two different scenario: Figure 8(a) presents the results with prior aggregation, while Fig. 8(b) without. Each run begins with the initialization of the CCU and ends when the decision process terminates. The two images demonstrate a reasonable memory footprint throughout the computation with an expected spike in correspondence with data decryption.
Refer to caption (a)
Refer to caption (b)
@@
(e) Encryption overhead To assess the overhead, we conducted experiments that account not only for the decision-making process itself but also for the data decryption phase. This step is essential in our system, as data must be decrypted after retrieval from IPFS and before decision execution (see the DECM step). We design two test configurations to evaluate the impact of different encryption strategies: light encryption, which serves as a baseline for comparison, and heavy encryption. Recalling the WS-ii step, in the heavy configuration, the CCU generates a different random number for each record, so it encrypts each such record with a different symmetric key (as per Tabs. II and II). For the light encryption variant, we modified the process described in WS-ii to use a single symmetric key (namely, we used the same random number for generating the symmetric key) for all entries. Indeed, this setup is strictly for experimental purposes. In practice, such an approach is discouraged, as compromising the single key would allow an attacker to decrypt the entire dataset rather than a single record. Figure 9 shows the execution time of the two encryption solutions. Figure 9(a) displays the execution time of light and heavy encryption strategies by varying the number of records with a fixed amount of columns ($28$), while in Fig. 9(b) we increase the number of columns with a fixed amount of users (16 00016\,000). The results highlight that the heavy encryption approach incurs a modest but reasonable overhead ($\approx\qty$80ms at most) compared to the light variant. This performance trade-off is justified because per-record encryption ensures significantly higher data confidentiality and resilience against key compromise.
Refer to caption (a)
Refer to caption (b)
@@
(f) TEE vs Server To quantify the overhead introduced by using both TEE and strong cryptographic guarantees, we compare our decision support system running inside the TEE with a standard server-based execution, inspired by [server1, server2]. Specifically, we modify our system to run outside the TEE but on the same machine. Figure 10 shows the comparison result: Figure 10(a) presents the execution time with increasing number of records and fixed columns, while in Fig. 10(b) we vary the number of columns with a fixed number of records (16 00016\,000). The results shows that the additional overhead is limited and primarily due to the decryption phase required by the TEE. In contrast, the server execution assume direct access to plaintext data within a trusted environment. Overall, our system provides strong security guarantees, including confidentiality for both data and decision logic, at a minimal performance cost and without compromising efficiency.
Refer to caption (a)
Refer to caption (b)
@@
Card Approval tests
@@
Figure 11: VAX, MED, and CARD datasets execution time
@@
Figure 11 summarizes execution times across all evaluated datasets. Notably, the MED dataset incurs higher decision evaluation time due to its larger rule set (see Sect. 6). However, decryption time remains consistent, underscoring SPARTA’s ability to maintain stable performance across domains with varying logic complexity and input scale.
7 Limitations and Future Work
Our solution presents some limitations. Firstly, our current design involves a single TEE. Even though SPARTA allows for the replication of CCUs, a load balancing strategy is not yet implemented. Such an extension would require non-trivial communication and session management adjustments, but would allow for a large-scale adoption. The memory limits of a TEE (see Sect. 2) present a scalability challenge when it comes to making decisions over a large volume of records (although the magnitude of affordable dedicated memory banks is reportedly increasing). Solutions to this problem can involve integrating Oblivious RAM (ORAM) techniques [ZeroTrace, Oblix, PRO-ORAM, Bulkor] or the execution of decisions in batched segments that can circumvent those memory constraints. In our threat model, we did not consider side-channel attacks that can compromise our solution’s security [SgxPectre, Cauldron, Inferring]. Although running SPARTA in a protected setting prevents such threats, several solutions exist to mitigate this type of attacks [Cipherfix, Strong, Obelix, HybCache]. We developed our solution for Intel SGX TEEs. Other solutions exist (e.g., AMD SEV, Arm CCA) and studying their employment is an interesting endeavor. Also, we will consider the integration of Intel Trust Domain Extensions (TDX). Our implementation does not support the entire FEEL language to write the filtering expressions and the aggregation functions. In particular, we currently back only one operation over the filtered data, and support only the standard aggregation functions (e.g., mean, sum, max, min). To further increase the expressive power of the decision making processes, we envision the adoption of TEE-based smart contracts for automated decision-making [cryptoeprint:2023/378, SECAUCTEE, Ekiden, SoK, ShadowEth]. Finally, we conducted experiments with public benchmarks showing good performance and scalability of our approach with low overhead. Nevertheless, an on-field use of SPARTA in real-world settings is paramount for validation purposes.
8 Related Work